Prepare the first row of a loan amortization schedule based on the following information. The loan amount is for $31,060 with an annual interest rate of 11.00%. The loan will be repaid over 28 years with monthly payments.

a) What is the Loan Payment?

b) What portion of this payment is Interest?

c) What portion of this payment is Principal?

d) What is the Loan balance after first monthly payment?
Answer #1
Note 1 i=interest rate per month=11/12=0.92%
n=time in months= 28*12=336
Note2 Monthly installment= Loan Amount
Pvaf(i,n)
= 31060
Pvaf(0.92,336)
= 31060
103.69
=                299.55
Note3 Loan amortisation table
Terms Opening Balance Interest installment paid Principal closing balance
1             31,060.00          285.75                  299.55                  13.80            31,046.20

a. Loan payment i.e. installment paid is $299.55

b.Interest portion is $ 285.75

c. Principal portion $13.80

d. Loan balance after 1st installment is $31046.2